OpenStudy (anonymous):

algebra help (solving for a and b) i am stuck on the algebra in a calculus problem 32= A(2k+1) + B(2k-1) 32= A2k + A + B2k - B 32= 2k(A+B) + (A-B) here is where im stuck i say 0 = A + B and 32 = A-B but its not right

OpenStudy (isaiah.feynman):

Look at the (2k-1) and (2k+1), do they remind you of anything?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

i distribute everything out and then im collecting like terms

OpenStudy (anonymous):

he might be talking about (x-1)(x+1)=(x^2-1) i believe but i dont think that will work

OpenStudy (anonymous):

@arturos your probably right, and your correct it doesnt help in this situation

OpenStudy (anonymous):

but i might be wrong

OpenStudy (anonymous):

what is k

OpenStudy (isaiah.feynman):

Only way I see here is to solve for A+B

OpenStudy (anonymous):

Isaiah is right you will probably have to isolate A then b with expanding

OpenStudy (anonymous):

a variable. \[32/((2k-1)(2k+1)) = A/(2k-1) + B/(2k+1)\]

OpenStudy (isaiah.feynman):

I don't understand. @clickspiker23

OpenStudy (anonymous):

WOW. ok so 0=A+B so B=-A which makes 32 = A-(-A0 32=2A A=16 B=-16

OpenStudy (anonymous):

which is the correct answer. Its been a long day

OpenStudy (anonymous):

yup all Isaiah

OpenStudy (isaiah.feynman):

Okay. That part was straight forward. I didn't look at it actually.

OpenStudy (anonymous):

im solving a telescoping series question

OpenStudy (anonymous):

thanks @Isaiah.Feynman and @arturos

OpenStudy (anonymous):

np